### v1

# Quantifying disorder through sharp spectroscopic fingerprints in an incompressible magnetic state

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Quantifying disorder through sharp spectroscopic fingerprints in an incompressible magnetic state
Chaebin Kim School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Sumedh Rathi School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Naipeng Zhang National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Arnab Seth School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Nikolai V. Simonov School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Aya Rutherford Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Long Chen Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Haidong Zhou Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Cheng Peng Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Mingyu Xu Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Weiwei Xie Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Advik D. Vira School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Mengkun Tian Institute for Matter and Systems, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Mykhaylo Ozerov National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Itamar Kimchi ikimchi3@gatech.edu School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Martin Mourigal mourigal@gatech.edu School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Dmitry Smirnov smirnov@magnet.fsu.edu National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Zhigang Jiang zhigang.jiang@physics.gatech.edu
School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA
(June 9, 2025)
Sharp spectroscopic fingerprints of disorder in an incompressible magnetic state
Chaebin Kim School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Sumedh Rathi School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Naipeng Zhang National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Arnab Seth School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Nikolai V. Simonov School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Aya Rutherford Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Long Chen Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Haidong Zhou Department of Physics and Astronomy, University of Tennessee, Knoxville, Tennessee 37996, USA    Cheng Peng Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Mingyu Xu Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Weiwei Xie Department of Chemistry, Michigan State University, East Lansing, Michigan 48864, USA    Advik D. Vira School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Mengkun Tian Institute for Matter and Systems, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Mykhaylo Ozerov National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Itamar Kimchi ikimchi3@gatech.edu School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Martin Mourigal mourigal@gatech.edu School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA    Dmitry Smirnov smirnov@magnet.fsu.edu National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA    Zhigang Jiang zhigang.jiang@physics.gatech.edu
School of Physics, Georgia Institute of Technology, Atlanta, Georgia 30332, USA
(June 9, 2025)
Disorder significantly impacts the electronic properties of conducting quantum materials by inducing electron localization and thus altering the local density of states and electric transport. In insulating quantum magnetic materials the effects of disorder are less understood and can drastically impact fluctuating spin states like quantum spin liquids. In the absence of transport tools, disorder is typically characterized using chemical methods or by semi-classical modeling of spin dynamics. This requires high magnetic fields that may not always be accessible. Here, we show that magnetization plateaus—incompressible states found in many quantum magnets—provide an exquisite platform to uncover otherwise undetectable amounts of disorder, regardless of the origin of the plateau. Using optical magneto-spectroscopy on the Ising-Heisenberg triangular-lattice antiferromagnet K2Co(SeO3)2 exhibiting a 1/3 magnetization plateau, we identify sharp spectroscopic lines, the fine structure of which serves as a hallmark signature of disorder. Through analytical and numerical modeling, we show that these fingerprints not only enable us to quantify minute amounts of disorder but also reveal its nature—as dilute vacancies. Remarkably, this model explains all details of the thermomagnetic response of our system, including the existence of multiple plateaus. Our findings provide a new approach to identifying disorder in quantum magnets.
Chemical disorder is a fundamental characteristic of solid-state materials and can manifest in various ways, such as impurities and dopants, site mixing and vacancies, bond disorder, and topological defects. In quantum materials, this disorder is imprinted at high temperatures during synthesis and becomes kinetically frozen, often remaining unnoticed until it influences putative emergent phenomena at low temperatures. However, disorder is not always detrimental in quantum systems. A notable example is the quantum Hall effect (QHE), where disorder plays a crucial role in the formation of quantized plateaus in the Hall resistance of devices and materials. This phenomenon has led to the establishment of a new international standard for electrical resistance [1, 2].
The role of disorder in quantum magnetic materials is less understood yet it is crucial for a wide range of systems, particularly quantum spin-liquids (QSLs). These systems are believed to host topological order and itinerant fractional excitations without exhibiting clear, conventional order-parameter-like signatures. The response, survival, and nature of potential QSLs in the presence of chemical disorder are not fully understood. Increasingly, researchers are recognizing that disorder can mimic the expected behavior of a QSL [3, 4, 5, 6, 7, 8]. But a significant challenge lies in quantifying the level of disorder using chemical techniques and translating this information into a model magnetic Hamiltonian, $\mathcal{H}$, whose emergent quantum many-body phases are not known in advance. A successful approach involves fitting inelastic neutron scattering or optical spectroscopy measurements under sufficiently high magnetic fields to polarize the system. In this semi-classical regime, it becomes possible to directly measure $\mathcal{H}$ [9], including the level of chemical disorder [10]. However, practical limitations arise from the need for large magnetic fields. Furthermore, spin-anisotropic systems of current interest may only reach saturation asymptotically [11].
The complex magnetization process of antiferromagnets may offer an alternative strategy. By utilizing quantized magnetization plateaus—when they occur—we show that it is not only possible to extract $\mathcal{H}$ at a much lower magnetic field, but these plateaus also exhibit unique, sharp spectroscopic signatures allowing for the detection of disorder with metrological precision. Similar to the QHE, this sensitivity emerges from the distinctive structure of incompressible magnetic states with quantized magnetization, an effect which has not been widely acknowledged until now.
Figure 1: Magnetization plateau and far-infrared magneto-spectroscopy measurements of K2Co(SeO3)2. a, Schematic representation of the normalized uniform magnetization $m=\langle S_{\rm tot}^{z}\rangle/NS$ of an ordered antiferromagnet described at the classical level (solid line) or including quantum corrections (dashed line). The spin-$S$ moments form a canted structure and become fully polarized at the saturation magnetic field $H_{\rm sat}$. b, Normalized uniform magnetization of an antiferromagnet with a field-induced quantized plateau state, which is stabilized between $H_{\rm c1}$ and $H_{\rm c2}$. The UUD state is a possible classical state realizing a $m=1/3$ magnetization plateau. c, Crystal structure of KCSO. Orange spheres represent oxygen, purple represents potassium, blue represents cobalt, and yellow represents selenium ions. The triangular lattice of magnetic Co2+ ions is represented by blue bonds, with the Ising-like nearest-neighbor exchange interactions represented by a brown ellipsoid. d, Frequency and magnetic field dependence of the normalized magneto-transmission spectra $I(\omega;H)$ of KCSO measured in Faraday geometry at $T\!=\!5$ K with the magnetic field direction $z$ along the $c$-axis, see Methods and Supplementary Information (SI) Sec. IA–B for details of our measurement and data analysis protocols. e, Energy modes $E(H)$ extracted from Lorentzian fits to the spectra at each magnetic field. Colors encode the nature of the underlying magnetic state: green and yellow for the compressible Y-like and V-like supersolid phases, and blue and light blue for the incompressible UUD phase corresponding to a $m=1/3$ plateau. Black lines are linear fits to the data using the averaged $g_{z}$ values, see SI Sec. IC for details. Dotted lines represent excitations originating from the Y and V phases, solid lines indicate dominant spin-flip modes in the UUD phase, and dashed lines correspond to disorder-induced satellite modes of the UUD phase.
At first glance, the magnetization process of antiferromagnets is simple. In the paramagnetic regime, Langevin or Brillouin theory describes the response of independent spins $S$—treated either as classical dipoles or quantum moments—to temperature $T$ and applied magnetic field $H$ [12]. In the magnetically ordered state, assuming that spins initially lie in the plane perpendicular to the magnetic field direction $z$, the classical dipoles cant toward the field direction. This induces a uniform magnetization $M(H)$ which increases linearly until $\langle S^{z}\rangle=S$ for each spin (Fig. 1a). The onset of this fully polarized state defines the saturation field $H_{\rm sat}$. A closer inspection reveals that non-linear contributions to the isothermal magnetization encode rich phenomena. For example, quantum corrections result in a convex magnetization in Heisenberg $S\!=\!1/2$ antiferromagnets [13, 14]. In geometrically frustrated systems, fractional magnetization plateaus can occur [15, 16, 17] with $\langle\sum_{i}S_{i}^{z}\rangle=\langle S^{z}_{\rm tot}\rangle\!=\!m\!\times\!NS$, where $m$ is a rational number and $N$ is the total number of spins (Fig. 1b). In systems with biquadratic exchange, quantum and classical effects compete in the stabilization of magnetization plateaus [18] while in frustrated $S\!>\!1/2$ systems with uniaxial anisotropy, a Devil’s staircase with an infinite number of magnetization plateaus is possible at the classical level [19]. Clearly, magnetization plateaus are a widespread and fascinating occurrence in magnetic systems [20]. They reflect that the total spin operator is conserved and commutes with the Hamiltonian, i.e., $[S^{z}_{\rm tot},\mathcal{H}]\!=\!0$. As the resulting magnetic excitations are gapped, a reliable determination of the system Hamiltonian $\mathcal{H}$ [21, 22] is possible at much lower magnetic fields than those necessary to reach a fully polarized state. Furthermore, regardless of the underlying classical or quantum mechanisms that stabilize the plateau, the resulting magnetic state is incompressible [20]; an increase in $H$ does not lead to an increase in $M(H)$. This incompressibility creates a highly sensitive platform for detecting even minute amounts of chemical disorder, because spins near a defect can respond strongly to the magnetic field, i.e. these spins are compressible.
To demonstrate this spectacular effect, we focus on the Heisenberg-Ising (or XXZ) triangular-lattice antiferromagnet K2Co(SeO3)2 (KCSO). This compound has long been known [23, 24] and was recently revisited in detail [25]. It features a robust magnetization plateau that spans an extraordinarily wide field range, from $2$ to $17$ T [26]. The magnetization plateau is flanked by compressible spin supersolid phases at low temperature [26, 27, 28]. These phases are the quantum analogs of non-collinear Y and V phases, where a static magnetic order in the $z$-direction coexists with phase coherence—but no static order—transverse to $z$. Although the exact nature of these phases is not yet fully understood, we will refer to them as Y and V in the rest of this article. The crystal structure of KCSO, depicted in Fig. 1c, comprises isolated $\rm CoO_{6}$ octahedra connected by $\rm SeO_{3}$ tetrahedra, forming a triangular lattice of $\rm Co^{2+}$ ions in the $ab$-plane. These triangular layers are stacked along the $c$-axis and separated by layers of K ions. Although the effective magnetic moments of KCSO extracted from the Curie-Weiss fit at high temperatures are nearly isotropic, with $\mu_{\rm eff}^{c}=6.14\mu_{\rm B}\approx 1.13\mu_{\rm eff}^{ab}$ (where $\mu_{\rm B}$ is the Bohr magneton) corresponding to an effective $g$-factor of $g_{c}=7.1$ [25], bilinear interactions exhibit a strong Ising character. The out-of-plane nearest-neighbor exchange $J_{zz}$ is an order of magnitude larger than the in-plane exchange $J_{xy}$ [26, 27]. Above the critical field $\mu_{0}H_{\rm c1}\approx 2$ T (where $\mu_{0}$ is the vacuum permeability) and at or below $T=5$ K, KCSO stabilizes an archetypal up-up-down (UUD) phase associated with a magnetization plateau at $m\!=\!1/3$. This state persists until the critical field $\mu_{0}H_{\rm c2}\approx 17$ T.
Figure 2: Schematic interpretation of magnetic excitations in the UUD phase of K2Co(SeO3)2. a, Magnetic excitations expected in the UUD phase with and without bond disorder. Solid lines (labeled #1 and #2) represent excitations in a clean system, while dashed lines (labeled #1a for above, and #1b, #2b for below) correspond to additional excitations in a disordered system. Blue and light blue indicate major spin-flip processes, while red and orange represent minor spin-flip processes. b, Spin structure of the ideal UUD phase for a $3\times 3$ triangular-lattice supercell. c, Visualization of major and minor spin-flip processes in clean and disordered systems. The flipped spin in each process is shown with a green sphere. The green bond represents a disorder bond with modified exchange interactions $J_{zz}\textquoteright=(1-d_{zz})J_{zz}$ and $J_{xy}\textquoteright=(1-d_{xy})J_{xy}$.
Here, we utilize far-infrared magneto-optical spectroscopy (FIRMS) and thermomagnetic measurements to investigate the behavior of KCSO near and within the plateau at $m=1/3$ (see Methods). To our surprise, although our crystals are grown using published protocols and chemical disorder ruled out by advanced X-ray diffraction and electron microscopy methods (see SI Secs. IV and V), the measurements reveal sharp spectroscopic fingerprints that can be clearly linked to very low levels of chemical disorder. These fingerprints were not seen in previous measurements [27]. Our optical measurements in the UUD phase reveal two dominant modes accompanied by three satellite modes, as illustrated in Fig. 1d and 1e. These satellite modes cannot be explained by clean-system models across three simulation techniques: linear spin wave theory (LSWT), Landau-Lifshitz dynamics (LLD), and exact diagonalization (ED). Instead, we develop a disordered-system model that fully captures our experimental observations. In this framework, the energy shift between the dominant and satellite modes is proportional to the strength of exchange disorder on nearest-neighbor bonds. The large energy shifts in our data suggest that the disorder in our KCSO crystals is more akin to spin vacancies than bond disorder. A typically undetectable amount of vacancies—as little as 2.5%—is sufficient to produce the observed intensity of the satellite modes. Moreover, magnetization and heat capacity measurements reveal a secondary magnetization plateau state with $m\!<\!1/3$, which is successfully reproduced in classical Monte Carlo (MC) simulations using the vacancy model.
Our FIRMS results, measured in Faraday geometry ($H\parallel c\equiv z$) at $T\!=\!5$ K, are shown in Fig. 1d. At low magnetic fields ($30<\mu_{0}H<$ 3 T), the system is in the Y phase [27], and the spectra exhibit two broad field-dependent modes that cross at $\mu_{0}H\approx$ 3 T. Around 3 T, the system enters the UUD phase and displays much sharper modes, with the two strongest intersecting at $\mu_{0}H\!=\!9.75$ T. Three satellite modes, running parallel to the main modes, are also observed. Above 15 T, an additional mode emerges and broadens with increasing field; we associate this mode with the onset of the V phase [27]. We determine the energy of major modes at each magnetic field, $E(H)$, using Lorentzian fits, and plot the results on Fig. 1e, where the color encodes the parent Y, UUD, or V phase associated with each mode. Now, we address the origin of the two strongest modes in the UUD phase. The magnetic Hamiltonian of KCSO in that field configuration reads
$\mathcal{H}=$ $\sum_{\langle i,j\rangle}[J_{xy}(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y})+J_{zz}S_{i}^{z}S_{j}^{z}]$ (1)
$+\mu_{0}\mu_{\rm B}g_{z}H\sum_{i}S_{i}^{z},$
where the first sum runs over nearest neighbor bonds, $J_{xy}>0$ and $J_{zz}>0$ are antiferromagnetic interactions, and $S_{i,j}$ are spin-1/2 operators. The second term describes the usual Zeeman effect, with magnetic field $H$ in the $z$-direction, and $g_{z}$ representing the $z$-component of the effective $g$-factor for the Co2+ ions. In the classical Ising limit, the excited states of the UUD phase can be expressed as integer multiples of $J_{zz}$:
$E_{n}=nJ_{zz}\pm\mu_{0}\mu_{\rm B}g_{z}H,~{}~{}(n=0,1,2,3,…).$ (2)
Figure 3: Simulated FIRMS spectra from our model of K2Co(SeO3)2 with and without disorder. a-c, Frequency and magnetic field dependence of simulated $I(\omega;H)$ for a clean system simulated using LSWT, LLD, and ED, respectively. d-f, Similar simulation approaches but for a disordered system. Gray lines represent the linear fits to mode energies, identical to those reported in Fig. 1e. Gray areas indicate the regime where the ground state of the system is not UUD. g, Same data as in Fig. 1d, shown for better comparison with the simulation results.
Due to optical selection rules, only transitions with $\Delta S^{z}=\pm 1$ are visible in FIRMS. Therefore, our experiment primarily detects excited states involving single or triple spin-flip processes (see SI Sec. II for further discussion). The two strongest modes observed in the UUD phase can be naturally attributed to single spin-flip excitations. In Fig. 2a, solid lines represent the mode energies $E(H)$ expected for single spin-flips allowed by the selection rule. These processes can occur at two distinct spin sites: up spin (major spin) and down spin (minor spin). A minor spin flip costs onsite energy in terms of $J_{zz}$ only—specifically, $3J_{zz}$ for the triangular lattice—since the flipped spin cannot propagate to neighboring spin sites. In contrast, a major spin flip incurs no onsite energy change in the classical limit but is accompanied by an additional quantum correction of $23J_{xy}/$ 2, reflecting that the flipped spin can propagate to three neighboring up-spin sites (Fig. 2b,c). Linear fits to these two modes (see SI Tab. S2) yield $J_{xy}$ = 0.20(1) meV, $J_{zz}$ = 2.88(1) meV, and $g_{z}$ = 7.43 (positive slope) or $g_{z}$ = 7.50 (negative slope), in good agreement with previous neutron scattering and magnetization studies [25, 26, 27] but with a more accurate determination of the $g$-factor. Triple spin-flip excitations, forming both bound-state and a free particle continuum, are also observed (see SI Sec . IIA-B for details).
Our data in the UUD phase exhibit three salient satellite modes that do not correspond to either single or triple spin-flip processes. To account for these excitations, we develop a disorder-bond model in which exchange interactions are modified as:
$J_{zz}^{\prime}$ $=$ $J_{zz}(1-d_{zz})~{}~{}{\rm and}$ (3)
$J_{xy}^{\prime}$ $=$ $J_{xy}(1-d_{xy}),$ (4)
where $d_{zz}$ and $d_{xy}$ represent the disorder-induced reduction of the respective interaction. Notably, this model yields fine spectroscopic features that are the unmistakable hallmark of the disorder’s nature. Specifically, a major spin-flip excitation can split into two modes, depending on whether the disorder bond is connected to a major spin (mode #1a in Fig. 2) or a minor spin (mode #1b in Fig. 2). The excitation energies are shifted by ±$\pm J_{zz}d_{zz}/2$, respectively. Moreover, when the disorder bond is connected to a major spin, there is a quantum correction of $-\alpha J_{xy}d_{xy}$ (where $\alpha\simeq n_{dm}$ is the disorder density in the dilute limit), due to the propagation of the flipped spin to two neighboring up-spin sites with intact bonds. In the case of a minor spin-flip process, the disorder only introduces an energy shift of $-J_{zz}d_{zz}/2$, since the flipped spin cannot propagate to neighboring spin sites (see detailed discussion in SI Sec. IID). More generally, we hypothesize that the splitting and multiplicity of spin-flip excitations in incompressible magnetic states serve as a unique indicator of bond disorder.
Figure 4: Magnetization and specific heat of our crystals of K2Co(SeO3)2. a, Temperature dependence of the magnetization of KCSO in a $c$-axis magnetic field. b-c, Simulated isothermal magnetization curves using the classical MC method for clean and disordered systems, respectively. The color of each curve corresponds to the temperature shown on the color bar. d, Schematic mechanism stabilizing an additional magnetization plateau around $\mu_{0}H=2.8$ T. The top and bottom panels show the spin configurations at 2 T and 4 T, respectively. Red, blue, and white circles indicate up spins, down spins, and vacancies, respectively. Orange circles highlight spin-flip sites that result in the additional plateau. e, Field derivative of the magnetization. f-g, Temperature and magnetic field dependence of the specific heat divided by temperature for KCSO in a $c$-axis magnetic field. The colorplot combines constant-field ($T\geq 0.8$ K) and constant-temperature ($T<0.8$ K) measurements.
Using the analytical solution of the disorder-bond model, we perform linear fits to the observed satellite modes (see SI Sec. IC). We find that the $y$-intercepts of these three modes are consistent with the model prediction, as illustrated in Fig. 2a. Quantitatively, we obtain $d_{zz}\!=\!0.93$ from the energy shift between modes #1 and #1b, and $d_{zz}=0.96$ from the energy shift between modes #2 and #2b. The close-to-unity value for $d_{zz}$ points to the complete absence of a bond, rather than a mere modification of the exchange due to disorder. The simplest mechanism to generate missing bonds is through spin-site vacancies, i.e., missing $\rm Co^{2+}$ ions.
Based on the above analysis and exchange parameters, we simulated the FIRMS spectra with and without bond disorder. We employed three different approaches, as summarized in Fig. 3: LSWT, LLD (to account for finite temperature), and ED (to account for quantum corrections). Calculations for a clean system, Fig. 3a-c, show that two sharp single spin-flip modes entirely dominate in the $m\!=\!1/3$ UUD phase, with only weak continua of triple spin-flip excitations in ED at high energies (see SI Sec. IIA). Simulations including disorder paint a different picture, Fig. 3d-f. To account for disorder in LSWT and LLD, we employ a spin vacancy model with two non-adjacent vacancies in a $9\times 9$ supercell. In ED, due to system size constraints, we use a broken-bond model with $d_{zz}=1$ for two bonds in a $6\times 3$ supercell, which equivalently reproduces the same excitations as the vacancy model (see SI Sec. IIC). As the vacancy/disorder density is directly related to the intensity ratio between the dominant and satellite excitations, we change the system size at a fixed number of defects until simulations are in excellent agreement with the data (SI Sec . IIB). Based on this ratio, we find that as little as 2.5% of $\rm Co^{2+}$ vacancies is sufficient to successfully reproduce all aspects of our experimental data in the UUD phase.
We now examine the impact of disorder on thermomagnetic properties of our crystals of KCSO. The isothermal magnetization $M(H)$ measured to sub-Kelvin temperatures with the field along the $c$-axis is shown in Fig. 4a. In the clean case, the magnetization is expected to reach the $m=1/3$ plateau in a single step (see Fig. 1b for a cartoon), as observed by previous experimental studies [26, 27] (see also SI Sec. III). In our case, however, an additional plateau with $m<1/3$ is observed from $\mu_{0}H\approx 2.8$ T, and the main plateau never reaches exactly $m=1/3$. To understand this behavior, we performed classical MC simulations for both our clean and disordered models of KCSO. In the clean case, the magnetization saturates at $m=1/3$ near $\mu_{0}H\approx 4$ T (Fig. 4b). In contrast, for a spin vacancy density of 2.5%, the magnetization shows a plateau around $\mu_{0}H\approx 2.8$ T, followed by a jump to the primary plateau at 3.4 T (Fig. 4c). The magnetization on both plateaus falls below $m=1/3$. By tracking the spin configuration at each magnetic field in our MC simulations, we find that the additional plateau originates from spin flips near vacancy sites (Fig. 4d). This additional plateau serves as a thermodynamic signature of disorder, but lacks the mechanistic insights provided by our spectroscopic approach.
Heat capacity measurements at low temperatures, with the magnetic field applied along the $c$-axis, are presented in Fig. 4f-g. The field dependence of specific heat at $T=0.2$ K reveals that a phase with gapless excitations exists below $\mu_{0}H\approx 1$ T, which we assign to the Y-phase or its quantum analog. As the field increases, the specific heat is suppressed but reaches a finite value again between $\mu_{0}H\approx 2.1$ T and 3.5 T (Fig. 4f). This enhancement survives up to $T\!=\!0.5$ K and matches the peaks observed in the field derivative of magnetization (Fig. 4e). This correlation suggests that the low-temperature specific heat can detect the disorder-induced plateau state through the density of states of local spin-flip processes near vacancies, as these become gapless. Above $T\!=\!1$ K, the heat capacity shows only a broad peak that corresponds to the field-driven transition from the Y phase to the UUD phase. These findings emphasize the significance of examining the low-temperature regime when investigating the effects of disorder, even in systems that exhibit Ising-limit behavior.
The spectacular spectroscopic signatures of disorder we observe in the $m\!=\!1/3$ plateau phase are not unique to KCSO. Additional measurements of the sister compound $\rm Rb_{2}Co(SeO_{3})_{2}$ (See SI Sec. IE) reveal similar satellite modes. Notably, the 2.5% vacancy density extracted in this work appears below the detection limit of laboratory diffraction and microscopy techniques commonly used in the routine structural characterization of quantum materials. By combining the physics of an incompressible magnetization plateau state with the high sensitivity of magneto-optical measurements, we demonstrate a viable route to bridge this technological gap. This approach not only provides quantitative insights into the amount of disorder, but also informs about its mechanistic origin. Given that magnetization plateaus are a common occurrence in frustrated and quantum magnets, and that magneto-optical techniques are becoming more accessible in laboratories worldwide, our method has the potential to accelerate the search for and validation of emergent magnetic states, such as QSLs.
Methods
Sample preparation
Single crystals of K2Co(SeO3)2 were prepared following the procedure reported in Ref. [25]. Specifically, high-purity and commercially available K2Co3, Co3O4, and SeO2 were mixed in a molar ratio of $1.2\!:\!0.33\!:\!2.2$ and loaded into an Al2O3 crucible, which was sealed in a vacuumed quartz tube. The tube was heated to $T\!=\!600~{}^{\circ}$, maintained at that temperature for 8 hours, and then cooled to room temperature at a rate of $6~{}^{\circ}$/hour. The resulting flat pink crystals were collected in water. The crystallographic $c^{\ast}$ and $c$ axes are normal to the flat surface of the crystals, as confirmed by backscattering Laue X-ray diffraction. Single crystals of Rb2Co(SeO3)2 were prepared similarly.
Far-infrared magneto-spectroscopy
Flat samples of appropriate size were selected for far-infrared magneto-spectroscopy (FIRMS) measurements using a Bruker VERTEX 80v Fourier-transform infrared (FTIR) spectrometer at the National High Magnetic Field Laboratory (NHMFL, Tallahassee, Florida). The sample was mounted with GE varnish on a nonmagnetic aperture (1.5 mm in diameter) and attached to a top-loading probe. The probe was then inserted into an optical cryostat equipped with a superconducting magnet reaching up to $\mu_{0}H\!=\!17.5$ T. Infrared light from a mercury lamp was delivered to the sample via an evacuated light pipe and focused onto the sample surface using a parabolic cone.
To reduce light-induced heating and improve signal-to-noise ratio, a 700 cm-1 low-pass filter was placed in the light path just above the sample. Transmitted infrared light passing through the sample was collected by another pair of parabolic cones and detected by a Si bolometer operating at liquid helium temperature. A schematic of the resulting Faraday transmission configuration, with the light propagating along $H\!\parallel\!c$, is shown in Fig. S1. The lowest effective sample temperature during measurements was $T_{\rm min}=5$ K, as measured by a thermometer mounted on the sample holder.
During the measurements, three different beam splitters (Multilayer Mylar, Mylar 50 $\mu$m, and Mylar 125 $\mu$m) were used to cover the relevant energy range. The raw spectra from each (beam splitter) measurement are shown in Fig. S2.
Spin dynamics simulation
Simulations of the transmitted frequency-dependent FIRMS spectra $I(\omega)$ were performed using the Sunny.jl package [29] and the QuSpin.py package [30, 31] by calculating the diagonal components of the frequency-weighted zero-momentum-transfer dynamical spin structure factor perpendicular to the light transmission direction,
$I(\omega)=\omega[S^{xx}({\bf q}=0,\omega)+S^{yy}({\bf q}=0,\omega)],$ (5)
where $z\!\parallel\!H\!\parallel\!c$. The dynamical spin structure factor was simulated using several approaches; in either case, magnetic moments were treated as dipolar spins and the system’s dynamics was governed by the same parent quantum Hamiltonian $\mathcal{H}$, Eq. 1, and its known exchange parameters.
The zero-temperature response ($T\!=\!0$ K) was simulated using linear spin wave theory, by which spin-1/2 operators ${\bf S}_{i}$ are transformed into Holstein-Primakoff bosons ${b}_{i}$ and the resulting bosonic Hamiltonian, truncated at quadratic order, is diagonalized using known numerical approaches implemented in Sunny.jl and other programs [32, 33].
The finite-temperature response was obtained in Sunny.jl using classical Landau-Lifshitz dynamics (LLD). For these simulations the spin system is assumed to be a product state $|{\bf\Omega}\rangle=\otimes_{i}|{\bf\Omega}_{i}\rangle$ over SU(2) coherent states $|{\bf\Omega}_{i}\rangle$ representing dipolar spin operators $\hat{{\bf S}}_{i}$, which are replaced with classical vectors ${\bf\Omega}_{i}=\langle{\bf\Omega}_{i}|\hat{\bf S}_{i}|{\bf\Omega}_{i}\rangle$. The time-dependent dynamics at finite temperature is calculated using the stochastic Landau-Lifshitz-Gilbert (LLG) equation:
$\frac{d\mathbf{\Omega}}{dt}=-\mathbf{\Omega}\times\left[\mathbf{\xi}+\frac{d\mathcal{H}_{\rm cl}}{d\mathbf{\Omega}}-\lambda\left(\mathbf{\Omega}\times\frac{d\mathcal{H}_{\rm cl}}{d\mathbf{S}}\right)\right],$ (6)
where $\mathcal{H}_{\rm cl}=\langle{\bf\Omega}|\mathcal{H}|{\bf\Omega}\rangle$ in the large-$S$ limit, $\mathbf{\xi}$ is a temperature-dependent Gaussian white noise, and $\lambda=0.2$ sets the coupling strength between the system and the thermal bath. The simulations were performed using a $9\times 9$ super-cell of the conventional chemical unit cell of the crystal with periodic boundary conditions. To match the overall excitation bandwidth of our data, we set the temperature to $T=0.5$ K in the LLD simulations. The system was first thermalized over 5,000 Langevin time steps with a step size of $\Delta t=0.013$ meV-1. After thermalization, five spin configurations were sampled using the LLG equation, with each configuration separated by 3,000 Langevin time steps to ensure decorrelation.
Exact diagonalization calculations were performed using QuSpin.py [30, 31]. The frequency-weighted dynamical spin structure factor, Eq. 5, was calculated using a $6\times 3$ supercell with periodic boundary conditions, employing the Lanczos algorithm [34] and the continued fraction method [35]. The number of Lanczos iterations was set to 200. Due to computational limitations, quenched chemical disorder was modeled using weak bonds rather than directly introducing vacancies. A more detailed discussion can be found in the Supplementary Information Sec. IIC.
Classical Monte Carlo simulation
Field-dependent magnetization was calculated from the standard Monte Carlo Metropolis algorithm using Sunny.jl package [29]. The supercell size for the Monte Carlo simulation was set to $18\times 18\times 1$. For each magnetic field, we perform 20000 spin-flip steps for the annealing at a given temperature. After the annealing process, 20 samples of spin configurations are averaged after 100 spin flips for each sample.
Thermomagnetic measurements
Isothermal magnetization measurements were performed using a Quantum Design MPMS-3 SQUID magnetometer operated in DC mode with the 3He insert option (temperatures from $T=0.4$ K to 4 K). The magnetic field was applied along the $c$-axis. Heat capacity measurements were conducted using a Quantum Design Dynacool PPMS equipped the Dilution Refrigerator unit. Magnetic fields up to $\mu_{0}H=5$ T were also applied along the $c$-axis to map the temperature- and field-dependent specific heat $C(T,H)$ using the thermal relaxation method over a temperature range from $T=200$ mK to 4 K. Measurements above $T=800$ mK were performed at fixed magnetic fields with a high density of data points along the temperature axis, while measurements below $T=800$ mK were conducted at fixed temperatures with a higher density of data points along the magnetic field axis.
Single-crystal X-ray diffraction
Single-crystal X-ray diffraction (SCXRD) was performed on a specimen with dimensions $0.15\times 0.1\times 0.08$ mm3. The crystal was affixed to a nylon loop using Paratone oil and examined with a Rigaku XtaLAB Synergy, Dualflex, Hypix SCXRD diffractometer, operated at room temperature.
[…]
Scanning transmission electron microscopy
A focused ion beam from a Thermo Fisher Helios 5CX was used to lift out a specimen from a K2Co(SeO3)2 single crystal. The specimen was then polished to a thickness of $<50$ nm to achieve atomic resolution using a Hitachi HD-2700 scanning transmission electron microscope (STEM). The microscope was operated at $\Delta V=200$ keV with a convergence angle of 27 mrad and a spatial resolution of $\Delta r\approx 1.3$. Annular dark-field (ADF) images were collected and compared with the K2Co(SeO3)2 lattice structure to determine the zone axis. A selected ADF image is shown in Fig. S10.
Acknowledgements.
We thank N. Peter Armitage for insightful discussions. This work (spectroscopy, thermomagnetic measurements, data analysis, modeling, interpretation and theory) was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under grants DE-FG02-07ER46451 (SR, NZ, NVS, DS, and ZJ), DE-SC-0018660 (CK and MM), and Early Career Research Program DE-SC-0025478 (AS and IK). The crystal growth effort at UTK (AR, LC, and HZ) was supported by the U.S. Department of Energy grant DE-SC-0020254. The SCXRD measurements at MSU (CP, MX, and WX) were supported the U.S. Department of Energy grant DE-SC-0023648. STEM measurements were performed at the GT Institute for Matter and Systems, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation under grant NSF-ECCS-2025462. FIRMS measurements were performed at the NHMFL, which is supported by the National Science Foundation Cooperative Agreement under grant NSF-DMR-2128556 and the State of Florida.
Author Contributions
SR, NZ, NVS., MO, and DS performed the FIRMS measurements. CK, HZ, and MM carried out the thermodynamic measurements. CK, SR, MM, and ZJ analyzed the experimental data, while CK performed the spin dynamics simulations under the guidance of IK, MM, and ZJ. AS and IK proposed the disorder-bond model. AR, LC, and HZ synthesized the single crystals. CP, MX, and WX performed the SCXRD measurements and fitting. ADV and MT conducted the STEM measurements. CK, SR, MM, and ZJ wrote the manuscript with input from all coauthors. ZJ and DS supervised the project.
Competing interests
The authors declare no competing interests.
Data availability
The data that support the findings of this study will be available during peer review in the Georgia Tech public data repository at: https://hdl.handle.net/1853/72971.
Codes availability
The codes that support the findings of this study are available from the corresponding authors upon reasonable request.
Supplementary Information
I Far-infrared magneto-spectroscopy measurements and data analysis
I.1 Experimental Setup
The far-infrared magneto-spectroscopy (FIRMS) measurement setup in the Faraday transmission configuration is illustrated (not to scale) in Fig. S1. To achieve the best signal-to-noise ratio in FIRMS measurements for the spectral range of interest in this work, we employed three beam splitters: Multilayer Mylar, Mylar 50 µm, and Mylar 125 µm. The spectral ranges of these beam splitters are specified in Tab. S1.
Figure S1: Schematic of the FIRMS measurement in the Faraday transmission configuration.
Table S1: Spectral ranges of the three beam splitters.
Manufacturer provided Ranges used in this work
Beam splitter Spectral range (cm-1) Spectral range (cm-1)
Multilayer Mylar 650–30 80–48
Mylar 50 µm 55–20 48–15
Mylar 125 µm 20–5 15–10
I.2 K2Co(SeO3)2 FIRMS Data Processing
The raw transmission spectra $\mathcal{T}(\omega)$ from FIRMS measurements are usually not directly useful for distinguishing magnetic excitations from the overall background, which includes absorption features from all objects in the light path. For the best presentation of the FIRMS data of K2Co(SeO3)2 (KCSO) measured at constant magnetic fields, we implemented different normalization methods, including:
  • Normalization by the zero-field spectrum: $(\omega;H)/\mathcal{T}(\omega;\mu_{0}H=0~{}\text{T})$ T )
  • Normalization by the maximum field spectrum: $(\omega;H)/\mathcal{T}(\omega;\mu_{0}H=17.5~{}\text{T})$ T )
  • Normalization by the averaged spectrum over all magnetic fields: $I(\omega;H)=\mathcal{T}(\omega;H)/\int\mathcal{T}(\omega;H^{\prime}){\rm d}H^{\prime}$
After visually comparing the three methods, we determined that the third method, normalization with the average spectrum, provided the best results. Accordingly, we use $I(\omega;H)$ throughout the remainder of this article.
Figure S2: (a-c) Normalized transmission spectra $I(\omega;H)$ of KCSO measured at different magnetic fields using (a) Mylar 125 µm, (b) Mylar 50 µm, and (c) multilayer Mylar beam splitters, respectively. All measurements were performed at $T=5$ K with a magnetic field increment of 0.25 T between successive spectra. (d) Combined data from the three different beam splitters.
Figure S2(a-c) shows the normalized transmission spectra, $I(\omega;H)$, of KCSO measured at different magnetic fields using three different beam splitters, while Fig. S2(d) presents the combined data, for the spectral range of each beam splitter summarized in Table S1. This approach eliminates noisy data caused by the low-energy cutoff or absorption lines of the beam splitters. Figures in the main text are generated based on Fig. S2(d).
I.3 Linear function fitting of extracted energy modes from FIRMS data
We extract the energy of nine distinct modes from Lorentzian fits to the normalized FIRMS spectra at each magnetic field (Fig. S3). Lorentzian profiles are expected to be an accurate method to determine peak energies given the high energy resolution and minimal peak distortion in our measurements. The extracted mode energies as a function of magnetic field read, $E(H)=E_{0}\pm\mu_{0}\mu_{\rm B}g_{z}H$, where the vacuum permeability $\mu_{0}$ and the Bohr magneton $\mu_{\rm B}$ are known paramaters. For each line, the fitted $y$-intercept $E_{0}$, and slope $g_{z}$ corresponding to the projection of the $g$-factor along the $z$ direction, are summarized in Tab. S2.
To understand the origin of each energy mode in the classical limit, it is convenient to evaluate the ratio of $E_{0}$ to the Ising interaction $J_{zz}$. In the Y and V phases, this ratio is close to an integer value, $E_{0}/J_{zz}\approx$ 2, 1, and $-1$ for all modes, see e.g. modes $\alpha$, $\beta$, $\delta$, and $\zeta$ in Fig. S3. However, in the up-up-down (UUD) phase, only two modes (1 and 2 in Fig. S3) exhibit ratios close to integers ($E_{0}/J_{zz}\approx$ 0 and 3). The other modes (1a, 1b, and 2b) show non-integer ratios, suggesting a different physical origin.
Figure S3: Energy modes $E(H)$ extracted from $I(\omega;H)$ and corresponding linear fits from Tab. S2, each labeled accordingly. Color indicates the underlying magnetic ground state: green for the Y phase, blue and light blue for the UUD phase, and orange for the V phase. Dotted lines represent excitations from the Y and V phases, solid lines indicate dominant spin-flip modes in the UUD phase, and dashed lines correspond to disorder-induced modes in the UUD phase.
Table S2: Energy modes extracted from the FIRMS spectra of Fig. S2 and corresponding linear fits $E(H)=E_{0}\pm\mu_{0}\mu_{\rm B}g_{z}H$.
[…]
$\alpha$ 46.7(5) 5.79(7)  -7.7(3) 2.01
$\beta$ 23.6(1) 2.93(2) +8.0(10) 1.02
$\delta$ 23.6(1) 2.93(2)  -7.0(20) 1.02
UUD-phase $E_{0}$ (cm-1) $E_{0}$ (meV) $g_{z}$ $E_{0}/J_{zz}$
1 2.43(7) 0.301(8) +7.43(1) 0.10
1a 10.26(6) 1.273(7) +7.52(1) 0.44
1b -8.1(1) -1.01(1) +7.49(2) -0.35
2 69.77(7) 8.650(9)  -7.50(1) 3.00
2b 58.7(1) 7.27(1) -7.50(2) 2.52
V-phase $E_{0}$ (cm-1) $E_{0}$ (meV) $g_{z}$ $E_{0}/J_{zz}$
$\zeta$21.1(4) -2.62(5) 7.46(7) -0.91
I.4 Rb2Co(SeO3)2 FIRMS Measurement
To demonstrate the generality of our findings in the main text, we performed FIRMS measurements on Rb2Co(SeO3)2 (RCSO), which possesses a similar crystal structure and magnetic properties as KCSO [25], but with a slight elongation along $c$-axis and thus a relatively weaker nearest neighbor interaction. Figure S4 shows $I(\omega;H)$ for a RCSO crystal grown in a similar fashion as KCSO and measured in the Faraday configuration at $T=5$ K. The dominant and fine spectroscopic features associated with the UUD phase in RCSO are in close agreement with KCSO with a slight reduction of the main mode crossing field from $\mu_{0}H\approx$ 9.75 T (KCSO) to $\mu_{0}H\approx$ 9.25 T (RCSO).
Figure S4: Field- and energy-dependent normalized transmission $I(\omega;H)$ of RCSO in the Faraday configuration at 5 K. The intensity scale is adjusted to highlight the fine spectroscopic features of the data.
II Modeling and interpretation
II.1 Classical picture of localized magnetic excitations in UUD phase
In the Ising limit, the energy of magnetic excitations in the UUD phase can be expressed classically as an integer multiple ($n=$ 0, 1, 2, 3, …) of the nearest-neighbor interaction $J_{zz}$, given by:
$E_{n,\pm}(H)=nJ_{zz}\pm\mu_{0}\mu_{\rm B}g_{z}H,$ (S1)
where the magnetic field $\mu_{0}H$ is applied along the $z$ direction. Due to the spin selection rule for optical transitions, our experiment can only detect spin-flip excitations corresponding to $\Delta S=\pm 1$. Figure S5(a,b) shows $E(H)$ diagrams and real-space pictures of the experimentally accessible excitations and their corresponding spin configurations. The solid lines in Fig. S5(a) represent the two possible excitations involving a single site (single spin-flip), while the dashed lines indicate the three excitations involving three sites and $\Delta S_{\rm tot}^{z}=\pm 1$ (triple spin-flip). In the main text, we do not emphasize the presence of two weak modes crossing at high energy around $E=62$ cm-1, see Fig. S5(c), which we assign to two out of three possible triple spin-flip excitations (modes #3 and #5). Since triple spin-flip excitations are higher-order excitations, their intensity is expected to be weaker than that of single spin-flips. This is confirmed by exact diagonalization (ED) simulations, as shown in Fig. S5(d). Note that in ED, the triple spin-flip excitations appear as a continuum as these simulations are performed for a finite transverse term $J_{xy}$ which leads to spin-flip hopping. The third triple spin-flip excitation (mode #4 in Fig. S5(a)) is not observed in either the experimental data or the ED simulations.
Figure S5: (a) Mode energies $E_{n,\pm}(H)$ of single spin-flip (solid lines) and triple spin-flip (dashed lines) excitations associated with the UUD phase. (b) Real-space visualization of the spin configuration for each mode, where the green sphere indicates the flipped spin(s). (c) Normalized transmission spectra $I(\omega;H)$ of KCSO measured in the Faraday configuration at $T=5$ K, with the two experimentally observed triple spin-flip modes highlighted. Circles denote excitation energies extracted from Lorentzian fits to the measured spectra. (d) Same ED simulation plots as in Fig. 3, with rescaled color contrast to highlight the presence of continua of triple spin-flip excitations.
II.2 Vacancy-density dependence of FIRMS intensity from spin dynamics simulations
To evaluate the vacancy density in our sample, we simulated its effect on the FIRMS intensity using linear spin wave theory (LSWT). Figure S6 shows the vacancy-density dependence of the transverse dynamical spin structure factor for ($6\times 6$), ($6\times 9$), ($9\times 9$), and ($12\times 9$) supercells. For each supercell, two non-adjacent vacancies were placed in the same configuration to reproduce a consistent spectrum, and the magnetic moment (and thus six bonds) associated with each of these vacancies was removed. Supercell sizes are multiples of 3 to accomodate the UUD magnetic structure. As expected, by comparing the integrated intensity ratios between the main peaks (modes #1 and #2 in Fig. S3) and the vacancy-induced satellite peaks (modes #1a-1b and #2b), we find that the ratio increases as the vacancy density decreases. From this, we determine that a vacancy density of no more than 2.5%, corresponding to two vacancies in the ($9\times 9$) supercell, best reproduces the intensity ratio observed in the experiment. The results are summarized in Tab. S3.
Figure S6: (Left) Vacancy-density dependence of the transverse dynamical spin structure factor calculated using LSWT. Gray areas indicate the regime where the ground state of the system is not UUD. (Right) Visualization of each supercell with spins and vacancies in the UUD phase. Red and blue represent up spin and down spin, respectively, while white denotes vacancies.
Table S3: Comparison of the average FIRMS intensity ratio between the main peaks (modes #1 and #2) and the satellite peaks (modes #1a-1b and #2b) from LSWT simulations and experimental data (for both KCSO and RCSO). The vacancy density corresponding to each supercell is indicated in parentheses.
Ratio $6\times 6$ (5.6%) $6\times 9$ (3.7%) $9\times 9$ (2.5%) $12\times 9$ (1.9%) K2Co(SeO3)2 Rb2Co(SeO3)2
$I_{1}/I_{1a}$ 3.11 5.40 10.37 13.52 8.14 3.87
$I_{1}/I_{1b}$ 2.34 3.93 7.82 10.43 7.13 12.84
$I_{2}/I_{2b}$ 2.49 4.00 7.38 31.13 9.3 5.72
II.3 Correspondence between vacancy model and broken-bond model
Due to memory-related computational limitations, we employed a broken-bond disorder model for ED simulations, by which disorder is introduced by removing individual bonds, while using the vacancy disorder model with a larger supercell for LSWT and classical dynamics simulations. In this section, we demonstrate that the two models are physically equivalent in describing magnetic excitations in the dilute vacancy density limit.
In the broken-bond disorder model, the exchange interaction of a broken bond is given by:
$J_{zz}^{\prime}$ $=$ $J_{zz}(1-d_{zz})~{}~{}{\rm and}$ (S2)
$J_{xy}^{\prime}$ $=$ $J_{xy}(1-d_{xy}),$ (S3)
where $d_{xy}$ and $d_{zz}$ represent disorder-induced reduction. In the vacancy disorder model, we set $d_{xy}=d_{zz}=1$ for the six bonds connecting the vacancy to its nearest neighbors. In the broken-bond model, we place two broken bonds in a ($6\times 3$) supercell. The ratio of broken bonds (i.e., the number of broken bonds to the total number of bonds) matches that of the vacancy model with two vacancies in a ($9\times 9$) supercell. As shown in Fig. S7, the resulting two-broken-bond model captures all three disorder environments for single spin-flip excitations present in the vacancy model and corresponding to modes #1a, #1b, and #2 in the experiment.
Figure S7: Visualization of the UUD spin configuration of the vacancy model using a ($6\times 6$) supercell (top) and the broken-bond model using a ($6\times 3$) supercell (bottom). Green bonds in the bottom panel represent broken bonds with $J_{xy}^{\prime}=J_{zz}^{\prime}=0$ or $d_{xy}=d_{zz}=1$. Light blue and orange circles indicate the single spin-flip processes involving a disordered environment.
II.4 Quantum corrections to the energy of the spin-flip excitations and their satellite modes
The quantum correction to the excitations of the classical Ising–$J_{zz}$ interaction is governed by:
[…]
where $S_{i,j}$ are spin-1/2 operators, and $S^{\pm}=S^{x}\pm iS^{y}$. As a generic effect of this term, the localized spin-flip excitations of both the clean and disordered model are allowed to propagate, hence the flat bands of these excitations split and gain bandwidth. We study this perturbatively to the first order of $J_{xy}$.
In the clean case, there are two kinds of single spin-flip excitations: major/up and minor/down spin flips, whose classical energy cost (zeroth order in $J_{xy}$) are given by 0 and $3J_{zz}$, respectively. To the first order in perturbation in $J_{xy}$, the major spin flips hop to the nearby major spin sites, hence giving rise to a hopping model on a honeycomb network formed by the up spins:
[…]
Here, $|i\rangle=S_{i}^{-}|\text{UUD}\rangle$ denotes a major spin flip at site $i$. The above can be solved via Fourier transformation to momentum space, giving rise to dispersion of $\varepsilon(\vec{k})=\pm{J_{xy}}|1+e^{i\vec{k}\cdot\vec{d}_{1}}+e^{i\vec{k}\cdot\vec{d}_{2}}|/{2}$, where $\vec{d}_{1,2}$ are the lattice vectors of the honeycomb network formed by the major spins (equivalently, the vectors connecting two second-neighbor major sites). The bandwidth of this branch (measured from the bottom of the lower band to maximum of the higher band) is estimated to be $3J_{xy}$. Since our optical measurements only see the transitions at $k=0$, major spin flip will have an energy correction of $23J_{xy}/$ 2. Note that although the first-order perturbation allows both signs (±$(\pm 3J_{xy}/2)$, we can only see the positive sign because of the symmetric wavefunction. On the contrary, the minor spin flips cannot propagate to the first order in $J_{xy}$, as the matrix element of $\mathcal{H}_{Q}$ is zero in the degenerate subspace of single minor flips.
Similar perturbative analysis can be applied to the disordered case to explain the observed satellite modes. As described in the main text, there is a single satellite mode originating from the minor spin flips with classical energy $23J_{zz}-J_{zz}d_{zz}/$ 2 and two other satellite modes from the major spin flips with energy $J_{zz}d_{zz}/2$ and $-J_{zz}d_{zz}/2$, depending on whether the disorder bond connects the flipped spin to major or minor spins.
The satellite mode of the minor flips does not get any correction due to $J_{xy}$ term as the quantum fluctuation cannot move such a spin flip even in the disorder-bond model. Among the satellite branches of the major spin flips, $J_{xy}$ term gives correction only to the higher energy mode because only in that scenario, the broken bond lies on the hopping path of the major flips. To estimate the correction, we consider a periodic arrangement of such bond disorder in the system such that each unit cell consisting of $\ell^{2}$ major spin sites contains a single broken bond. Hence, the disorder density is given by
$n_{d,m}=\frac{\text{Number of broken bonds}}{\text{Number of major sites}}=\ell^{-2}=\frac{9}{2}\frac{\text{Number of vacancy bonds}}{\text{Number of total bonds}}=\frac{9n_{d}}{2}.$ (S6)
The dilute disorder density limit which is relevant for this work is obtained at large $\ell$ limit. The advantage of this periodic arrangement is that it restores the translation symmetry even in the disordered system, hence we can compute the band structure of such tiling. The band minima of the upper satellite mode of the major spin-flip excitations then modifies as:
$\frac{J_{zz}d_{zz}}{2}+\frac{3J_{xy}}{2}\rightarrow\frac{J_{zz}d_{zz}}{2}+J_{xy}\left(\frac{3}{2}-\alpha d_{xy}\right),$ (S7)
where $\alpha$ generically depends on the disorder density. We plot $\alpha$ in Fig. S8 as a function of $n_{d,m}$ for various values of $d_{xy}$. For small $d_{xy}$, $\alpha$ scales linearly with disorder density, $\alpha=c_{0}n_{d,m}$ with $c_{0}\approx 1$. For larger $d_{xy}$, which is relevant for the broken-bond model, the scaling starts slowly deviating from the linear behavior. For $d_{xy}=1$, we obtain the fit to be $\alpha=0.9n_{d,m}^{1.13}$. We note that the accurate value of coefficient of the power law may further vary depending on the detailed structure of the defect, e.g., whether we consider broken-bond or vacancy model. But its generic dependence on the disorder density remains robust at the dilute $n_{d,m}$ limit.
Figure S8: Quantum correction to the upper satellite branch of the major spin flips (log-log scale). Different colored data points represent different disorder strength ($d_{xy}$). For small $d_{xy}$, $\alpha\approx n_{d,m}$ becomes a good fit, which is shown in black dashed line. As the disorder strength increases, the power law starts deviating from 1. For the broken-bond case ($d_{xy}=1$), the fit is shown by the gray dashed line.
III Sample dependence of magnetization data
To assess the sample dependence of vacancy density in KCSO, we compared our low-temperature magnetization data ($T<0.5$ K) with previous studies [26, 27]. The saturated plateau at 7 T in our data clearly indicates that our sample does not reach the ideal 1/3 magnetization value reported in earlier works (Fig. S9). Furthermore, previous studies did not observe an additional pre-plateau phase in the low-field region $\mu_{0}H<3.5$ T, further suggesting the presence of vacancies in our sample.
Figure S9: Comparison of our magnetization data with previous studies. Our results and the reference data from Ref. [27] were measured at 0.4 K using MPMS-3 and 0.5 K using PPMS, respectively. The reference data from Ref. [26] were measured at 0.19 K using Faraday balance magnetometry.
IV Detailed fitting results of single-crystal X-ray diffraction (SCXRD)
[…]
Chemical Formula K2Co(SeO3)2
Formula weight 391.05 g/mol
Space group $R\bar{3}m$
$a=5.5225(2)$
Unit cell dimensions $b=5.5225(2)$
$c=18.5390(8)$
Volume 489.66(4) Å3
$Z$ 3
Density (calculated) 3.978 g/cm3
Absorption coefficient 18.245 mm-1
F (000) 543.0
[…]
Reflections collected 4855
[…]
Refinement method Full-matrix least-squares on F2
Data/restraints/parameters 441/0/16
[…]
$R_{1}$ (all) = 0.0211; $wR_{2}$ (all) = 0.0455
Largest diff. peak and hole $+1.05$ e-3 and $-0.92$ e-3
RMS deviation from mean 0.188 e-3
Goodness-of-fit on F2 1.106
[…]
Table S5: Atomic coordinates and equivalent isotropic atomic displacement parameters (Å2) of KCSO with fixed occupancy. (Ueq is defined as one-third of the trace of the orthogonalized Uij tensor.)
K2Co(SeO3)2 Wyck. $x$ $y$ $z$ Occ. Ueq
Se1 $6c$ 0.3333 0.6667 0.79781(2) 1 0.00833(9)
Co1 $3b$ 0.6667 0.3333 0.8333 1 0.00887(10)
K1 $6c$ 0.6667 0.3333 0.63940(3) 1 0.01892(12)
O1 $18h$ 0.49424(11) 0.50576(11) 0.75946(7) 1 0.0144(2)
Table S6: Atomic coordinates and equivalent isotropic atomic displacement parameters (Å2) of KCSO with the Co site occupancy treated as a fitting parameter. (Ueq is defined as one-third of the trace of the orthogonalized Uij tensor.)
K2Co(SeO3)2 Wyck. $x$ $y$ $z$ Occ. Ueq
[…]
Table S6 shows the analysis results with the Co site occupancy treated as a fitting parameter. No significant differences are observed between Tables S6 and S6, indicating that the vacancy density in KCSO is below the detection sensitivity of SCXRD.
V Scanning transmission electron microscopy (STEM)
To visualize the crystalline structure and identify potential defects within our samples, STEM was performed. Figure S10 shows the annular dark-field (ADF) image of KCSO along the [110] direction, overlaid with the crystal structure. The alignment between atom positions (Co in blue, K in purple, and Se in green) and ADF bright spots demonstrates good agreement. No visible defects were observed.
Figure S10: ADF image of a single-crystal KCSO along the [110] direction, overlaid with the crystal structure.